# Practitioners’ Discussions on Building LLM-based Applications for Production

### v1

@@ Abstract
Background: Large language models (LLMs) have become a paramount interest of researchers and practitioners alike, yet a comprehensive overview of key considerations for those developing LLM-based systems is lacking. Objective: This study addresses this gap by collecting and mapping the topics practitioners discuss online, offering practical insights into where priorities lie in developing LLM-based applications.
Method: We collected 189 videos from 2022 to 2024 from practitioners actively developing such systems and discussing various aspects they encounter during development and deployment of LLMs in production. We analyzed the transcripts using BERTopic, then manually sorted and merged the generated topics into themes, leading to a total of 20 topics in 8 themes.
@@
Implications: Our results highlight current discussions and challenges in deploying LLMs in production. This way, we provide a systematic overview of key aspects practitioners should be aware of when developing LLM-based applications. We further pale off topics of interest for academics where further research is needed.
@@ I Introduction
The rapid evolution of Large Language Models (LLMs) in recent years has advanced the state of the art across multiple software engineering research fields, including code generation [1, 2], automated testing [3, 4], and comment generation [5]. Despite the successful application of LLMs in numerous research tasks, little work has been done on how LLM applications are built and deployed in practice. Although experience reports exist [6] and a systematic study addresses challenges in building LLM-based components [7], there is still no comprehensive overview of what practitioners consider important across the entire lifecycle of LLM applications.
@@
We know from traditional ML-enabled software system that productionization comes with its own challenges. Interviews, experience reports, and user studies draw a diverse picture of themes to consider. For instance, socio-technical challenges due to scattered organizations or lack of data science skills in the management cause productionization failures [8], communication challenges between software engineers and data scientists hinder the transfer of ML models to developers [9], and handling, managing, and versioning the data introduces a whole new dimension in the software life-cycle [10]. So, already the development of traditional ML-enabled software system is unique in many facets, it is, thus, crucial to obtain such a deep understanding when it comes to developing LLM-based applications.
@@
LLMs distinguish themselves from traditional ML models through several key aspects that alter the conventional machine learning workflow. Unlike traditional approaches, which necessitate the collection, processing, and domain-specific training of data, LLMs undergo a one-time pre-training process and are subsequently adapted to specific domains via methods, such as prompt engineering and fine-tuning. Moreover, the development of LLMs typically demands extensive teams with specialized expertise and significant hardware resources, contrasting with the smaller, in-house efforts typical for traditional ML models. Consequently, the resource-intensive nature of LLMs compels many companies to depend on external services for deployment and provisioning, adopting an LLM-as-a-service model, which might hold its own challenges and best practices. The largest difference, however, may lay into the black-box nature of LLMs that makes them notorious challenging to trust or even debug their output. This aspect alone calls for many techniques and methodologies to use an LLM in a practical setting.
@@
Given these fundamental differences, we cannot necessarily rely on existing knowledge of engineering challenges for ML-enabled systems. Thus, the goal of this work is to provide an overview of current topics and challenges practitioners discuss online and highlight key aspects that should be considered when building LLM-based applications. To accomplish this, we conduct a semi-automated thematic analysis of 92 hours of talks and conversations recorded by practitioners, which have been made publicly available on YouTube. Analyzing these public videos has been shown to contain valuable insights, that can be of comparable depth as traditional interview studies [8]. Not only that, they also provide an unfiltered view on what practitioners think is worth discussing with the community.
@@
Our results show that the discussed topics fall within eight main themes Architecture & Design, Model Capabilities & Techniques, Tools & Infrastructure, Evaluation, Risks & Ethics, Monitoring, Costs, and Output Verification. Here, we can see that some themes stem from the incorporation into a software system (e.g., Architecture & Design), some already known from ML-enabled software systems (e.g., Tools & Infrastructure and Monitoring), and some are entirely new (e.g., Costs and Output Verification). Within these themes, we find 20 distinct topics, of which the most prevalent one is concerned with retrieval augmented generation (RAG) systems.
@@
  • A comprehensive replication package, including a mapping of videos to relevant topics [11].
@@ II-A Challenges in building AI-enabled systems
While prior studies extensively explored challenges in building ML-enabled software systems, our work focuses specifically on the themes and topics unique to developing LLM-based software. Although LLMs are part of the broader machine learning landscape, they have only recently become accessible to companies of all sizes. Additionally, LLMs have unique properties that distinguish them from traditional ML models, such as requiring vast datasets and computational resources, exhibiting few-shot and zero-shot capabilities, and often relying on prompt engineering to adjust behavior. These factors impose unique challenges and areas of interest for practitioners. Therefore, our study focuses on topics relevant to practitioners by analyzing online discussions to provide a comprehensive overview of considerations for building LLM-based applications.
@@ II-B Challenges in building LLM-based applications
Figure 1: Overview of the different steps of our applied methodology.
While these studies offer preliminary insights into the challenges and solutions practitioners face when building LLM-based applications, their scope remains focused on specific areas. By contrast, our study aims to provide a broader overview of themes generally relevant to practitioners, covering areas such as architecture, costs, infrastructure, and model capabilities. With this work, we complement and extend the existing body of research and present a thematic map of practitioners’ interests that can guide companies when starting LLM initiatives and help researchers identify yet underexplored topics with high industrial relevance.
@@ III Methodology
We repeated this process three times in 11 months. Repeating this process at three different time points allowed us to capture a more representative set of videos, as content on YouTube is continuously updated. This approach ensured that newer, potentially relevant videos published after the initial searches were included, improving the timeliness of our dataset.
@@
We transcribed the filtered videos using whisper-large [23], generating one transcript per video. Next, we applied a semantic text splitter [24], to obtain disjunct, contextually coherent paragraphs of 300 tokens maximum, which corresponds to approximately 225 words, on average. This way, we yield 4100 paragraphs on which we applied topic modeling using BERTopic [25].
@@ IV Results
Next, we give an overview of our identified themes, including a discussions on the individual aspects of each topic, for which we provide direct quotes. Each quote is set in gray color and includes a subscript identifier, linking it to the corresponding video in the references. We further investigate co-occurring topics in videos.
@@ IV-A Themes of Building LLM-based Applications
The most frequently mentioned topic across all themes is retrieval-augmented generation (RAG) [26]. While LLMs encode vast amounts of information, they are limited to the data available during training, missing out on newer information or proprietary, sensitive data. RAG addresses this limitation by leveraging a vector database to store and retrieve prior ingested data. Before generating a response, the database is queried for relevant documents that are subsequently added to the prompt. This does not only provide timely data, but also reduces hallucination by providing additional context to the prompt.
@@
Speakers explain some of the optimization challenges in greater detail, such as difficulties and trade-offs when determining a good chunk size. Chunking the data is usually done before ingestion in the database, when larger documents likely would not fit in the context window of the LLM or only parts of the document contain relevant information. However, chunking long documents might also lead to adverse cuts in context. One speaker explains this with an example of an article about Alan Turing, which would likely be not split correctly, such that the retrieval will [..] get me like the first three chunks, but the last three chunks, which also corresponds to the latter half of the document may not have a mention of Alan TuringV66. So, determining a good chunk size is crucial. One speaker advocates for a human-in-the-loop process to determine good chunk sizes for different use cases: you’ll have a human who’s validating [..] this chunk size for this type of document is giving the best resultsV147. Newer approaches, such as contextual retrieval embed document context within individual chunks, but at the cost of larger chunk sizes.
The data within the vector database is stored as vectors, created by embedding models. How embeddings are created an updated is a frequent topic since the quality of your embeddings really drive the quality of a retrieval modelV15. However, pretrained embedding models might deliver an inadequate representation of your data [..].V32 and fine-tuning the embedding model is suited because [..] shifting the embedding points does help to improve performanceV32. Additionally, there is often the need to frequently update and extend data within the database to ensure there is no outdated information, which is a challenge on its own. Nonetheless the bigger problem is if the embedding models have to be exchanged or fine-tuned to account for new data, it may require re-indexing the whole database. This can become [..] very expensive and slow.V32. A potential solution lies in not fine-tuning the embedding model [..] but [..] fine tune some sort of adapter on the query and just keep the document embeddings [..] frozenV32. So, while RAG systems are seen as promising architectures, they pose their own challenges and trade-offs that need to be considered.
@@
A key factor for improving vanilla LLMs is to enhance the context length of the query, allowing the LLM to generate responses based on a larger quantity of information. However, the token size is usually fixed and it is up to the engineer to decide, which information to store for later inclusion and which to drop. Thus, memory management within applications is essential to handle context efficiently. One speaker describes this as: At the end of the day, when I talk about memory in all our applications, what I really mean is remembering previous interactions and then using those to inform future interactionsV174.
Effective memory management involves selecting only the most relevant information to retain, rather than preserving all historical data, which ensures models remain efficient and avoid confusion from outdated or irrelevant context. Techniques, such as conversation buffer memory and conversation summary memory have been proposed by the practitioners: the former keeps recent details accessible, while the latter provides a condensed summary of past interactions, balancing specificity and broader context. As another speaker notes, this method is an interesting way to address that problem for a chatbot-like scenarioV131 by focusing on the last $K$ relevant messages and managing context length effectively.
@@
A key insight shared was the importance of structured prompts, such as using XML tags or fixed prompt placement, to help models interpret context and improve performance. Additionally, certain prompt quirks—like using specific phrases or even statements like ”take a deep breath” or ”ensure the scores are correct”, sometimes enhance model responses, though the reasons remain unclear: So nobody knows why [..] you introduce like a statement, like ‘take a deep breath’ and the model tends to like do this.V66. This lack of theory makes prompting somewhat experimental, with practitioners trying unconventional statements based on observed improvements.
@@
Model size is not only a relevant for flexible deployment and low latency, but generally indicates the model’s capabilities and task performance. For many open ended tasks, a larger model size often leads to a higher quality of results, as it has been also shown via the observed scalability laws of LLMs [32]. Practitioners discuss this trade-off between model size and performance, alongside other relevant aspects, such as deployment efficiency. Currently, larger models exhibit better reasoning abilities but come with higher serving costs, requiring specific hardware and skills: [By using adaptation techniques] is how you’re able to build a much, much smaller model, which is equally as accurate, which is much, much easier to deploy.V96. Overall, this topic is closely related to specific techniques to improve model performance, such as fine-tuning or prompt engineering, as these technique generally hold the potential to tailor models towards a specific use case, and improve performance, especially for smaller model.
@@
In some cases practitioners discuss how to optimize LLMs such that serving them is easier and less costly, while retaining performance on specific tasks. The techniques mentioned within discussions are quantization, knowledge distillation, and engineering optimizations. Quantization [33], for instance, reduces the size of the stored model weights to lower memory requirements and computations at the cost of reduced accuracy of the model. Knowledge distillation [34] in contrast, relies on training a smaller model as a student to replicate the behavior of the bigger model. Most parts of this topic correspond to explanations of these techniques, alongside pros and cons.
@@
There is lot of discussion about whether the LLM should be deployed in a dedicated cloud, on-premise, or whether using an external LLM API, such as provided by OpenAI [^1] https://openai.com/. This decision is not easy: speakers discuss several considerations, such as latency, performance, costs, and control. Hosting models on-premise or in a dedicated cloud has several advantages such as more predictable long term costs and better compliance and control, as noted by one speaker, [hosting your own model is cheaper] since you can control it, it’s a lot more predictableV16. Conversely, choosing LLMs-as-a-service is more convenient, has lower upfront costs and higher scalability. However, speakers mention several challenges they encountered with LLM APIs, such as rate limits, timeouts, and unpredictable downtimes, exemplary indicated by one speaker: [some days] you’ll just have really bad latency and your HTTP calls will time outV19. Another speaker mentions that they rely on a hybrid approach, using an external LLM API not for the latency sensitive use casesV17, but rather batch jobs that are not time sensitive. Hence, where to put the LLM is an outreaching architectural decision that needs to be made on the tasks, resources, and environments at hand.
@@
Managing the infrastructure for training and inference of LLMs is a challenge on its own. One frequently mentioned tool is Kubernetes; a container orchestration platform that is already widely used for managing traditional ML workflows. One speaker, however, explicitly states that these workloads are not necessarily comparable [..] a lot of machine learning 1.0 workloads were batch and I would say more the 2.0 like generative AI, most of that’s actually streamingV11. This possesses unique challenges for the kind of infrastructure that is needed and the tools to manage it. Thus, other hosting and optimization frameworks are necessary. One speaker mentions that they finally used Ray to spread the model across different instancesV103 as it was not possible to fit a model on a single Kubernetes node. Key considerations discussed are cost optimization and performance balancing that both are influenced by different techniques, such as load balancing, dynamic batching, and efficient resource allocation. LLM-specific techniques and patterns may evolve that have the potential to improve reliability and performance, such as described by one speaker: You need some form of distributed queue that is aware of the length of the completion and the batch sizes and that schedule is based on that.V51.
@@
Use case specific evaluations with humans in the loop are often seen as gold standards, but are expensive and not scalable. An alternative is using LLMs to evaluate the output of another model, which is seen as very powerful as a nything you can prompt a model to do, you can get it to evaluate.V3. While promising, this method also comes with its own challenges, such as ensuring that the evaluating model’s biases do not influence the assessment. Alongside different evaluation settings, speakers mention that having the right data and metrics is key for successful evaluation that builds trust throughout development cycle, from prototyping to monitoring and regression testing.
@@
A key challenge of using LLMs in an application is dealing with hallucinations, that is, when the model generates output that is factually incorrect or misleading. One speaker mentions that hallucinations might occur due to out-of-distribution requests, but [..] that quantifying what’s out of distribution is a lot more difficult [than in traditional ML].V4. Practitioners propose several strategies to mitigate hallucinations, such as prompt chaining and self-critique. While the first one is about combining different prompts within one, the latter works by explicitly asking the model to reflect on its output: [..] you’re asking it to self critique, right? ‘Hey, does this make sense to you?’ And the model might say, ‘oh, I apologize. That actually doesn’t make sense’ and gets it right the second time.V3. Finally, one speaker indicates that by using LLMs, we can never fully exclude hallucinations and concludes: [..] well if my application cannot tolerate you know hallucinations maybe I shouldn’t use it for that application.V70.
@@
A recurring insight is that while larger context length is generally better since it allows LLMs to incorporate more in-context information and longer inputs, it is significantly more expensive as [..] when you take a bigger context window, [..] there is a quadratic increase in the cost of training.V159. However, it is also noted that smaller context windows enables quicker experimentation and prompt adjustment without excessive spending. So, these discussions emphasize the importance of finding an optimal context length, as it has substantial cost and performance implications for LLMs.
@@
Speakers discuss different tools, which work as an intermediary between the user and LLM and taking care of the interactions by providing several mechanisms for handling LLM responses, ensuring correctness, safety, and compliance. Besides verfiying outputs, these tools may include mechanisms to directly correct LLM output by automatically re-asking the model for correct or compliant output.

### v2

# Themes of Building LLM-based Applications for Production: A Practitioner’s View

@@ Abstract
Background: Large language models (LLMs) have become a paramount interest of researchers and practitioners alike, yet a comprehensive overview of key considerations for those developing LLM-based systems is lacking. This study addresses this gap by collecting and mapping the topics practitioners discuss online, offering practical insights into where priorities lie in developing LLM-based applications.
Method: We collected 189 videos from 2022 to 2024 by practitioners actively developing such systems and discussing various aspects they encounter during development and deployment of LLMs in production. We analyzed the transcripts using BERTopic, then manually sorted and merged the generated topics into themes, leading to a total of 20 topics in 8 themes.
@@
Implications: Our results highlight current discussions and challenges in deploying LLMs in production. This way, we provide a systematic overview of key aspects practitioners should be aware of when developing LLM-based applications. We further highlight topics of interest for academics where further research is needed.
Index Terms:
LLM in production, RAG, retrieval-augmented generation, fine-tuning, prompt engineering, agents, evaluation
@@ I Introduction
The rapid evolution of Large Language Models (LLMs) in recent years has advanced the state of the art across multiple software engineering research fields, including code generation [1, 2], automated testing [3, 4], and comment generation [5]. Despite the successful application of LLMs in numerous research tasks, little research has been done on how LLM applications are built and deployed in practice. Although experience reports exist [6] and a systematic study addresses challenges in building LLM-based components [7], there is still no comprehensive overview of what practitioners consider important when building and deploying LLM applications.
@@
We know from traditional ML-enabled software systems that productionization comes with its own challenges. Interviews, experience reports, and user studies draw a diverse picture of themes to consider. For instance, socio-technical challenges due to scattered organizations or lack of data science skills in the management cause productionization failures [8], communication challenges between software engineers and data scientists hinder the transfer of ML models to developers [9], and handling, managing, and versioning the data introduces a whole new dimension to the software life-cycle [10]. So, already the development of traditional ML-enabled software system is unique in many facets, it is, thus, crucial to obtain such a deep understanding when it comes to developing LLM-based applications.
@@
LLMs distinguish themselves from traditional ML models through several key aspects that introduce unique challenges to the conventional machine learning workflow. Unlike traditional models, which are trained on domain-specific data entirely, LLMs are pre-trained on extensive general-purpose data, leading to difficulties in aligning outputs to specific domains, unpredictable behavior, and challenges in integrating their black-box nature into trust-dependent applications. Moreover, the development of LLMs typically demands extensive teams with specialized expertise and significant hardware resources, particularly due to their scale and the computational demands of fine-tuning or serving pre-trained models, which sets them apart from smaller, domain-specific traditional ML efforts. Consequently, the resource-intensive nature of LLMs compels many companies to depend on external services for deployment and provisioning, adopting an LLM-as-a-service approach, which might hold its own challenges and best practices. However, the largest difference may be the black-box nature of LLMs that makes them notoriously challenging to trust and even debug. While neural networks in general exhibit black-box characteristics, LLMs amplify this challenge due to their latent behaviors, the opacity of their training data and processes, and the inability to explain outputs even in high-stakes applications. This aspect alone calls for novel methods to use LLMs in a practical setting.
@@
Given these fundamental differences, we cannot necessarily rely on existing knowledge of engineering challenges for ML-enabled systems. Thus, the goal of this work is to provide an overview of current topics and challenges practitioners discuss online and highlight key aspects that should be considered when building LLM-based applications. To accomplish this, we conducted a semi-automated thematic analysis of 92 hours of talks and conversations recorded by practitioners, which have been made publicly available on YouTube. Analyzing these public videos has been shown to contain valuable insights that can be of comparable depth as traditional interview studies [8]. Not only that, they also provide an unfiltered view on what practitioners think is worth discussing with the community.
@@
Our results show that the discussed topics fall within eight main themes Architecture & Design, Model Capabilities & Techniques, Tools & Infrastructure, Evaluation, Risks & Ethics, Monitoring, Costs, and Output Verification. Here, we can see that some themes stem from incorporation into a software system (e.g., Architecture & Design), some are already known from ML-enabled software systems (e.g., Tools & Infrastructure and Monitoring), and some topics are entirely new to this extent (e.g., Costs and Output Verification). Within these themes, we find 20 distinct topics, of which the most prevalent one is concerned with Retrieval Augmented Generation (RAG) Systems.
@@
  • A comprehensive replication package, including a mapping of videos to relevant topics  [^1] https://doi.org/10.5281/zenodo.14753685.
@@ II-A Challenges in building AI-enabled systems
While these studies extensively explored challenges in building ML-enabled software systems, our work focuses specifically on the themes and topics unique to developing LLM-based software. Although LLMs are part of the broader machine learning landscape, they have only recently become accessible to companies of all sizes. Additionally, LLMs have unique properties that distinguish them from traditional ML models, such as requiring vast datasets and computational resources, exhibiting few-shot and zero-shot capabilities, and often relying on prompt engineering to adjust behavior. These factors impose unique challenges and areas of interest for practitioners. Therefore, our study focuses on topics relevant to practitioners by analyzing online discussions to provide a comprehensive overview of considerations for building LLM-based applications.
@@ II-B Challenges in building LLM-based applications
Figure 1: Overview of the different steps of our applied methodology.
@@
While these studies offer preliminary insights into the challenges practitioners face and solutions they have when building LLM-based applications, their scope remains focused on specific areas. By contrast, our study aims to provide a broader overview of themes generally relevant to practitioners, covering areas such as architecture, costs, infrastructure, and model capabilities. With this work, we complement and extend the existing body of research and present a thematic map of practitioners’ interests that can guide companies when starting LLM initiatives and help researchers identify yet underexplored topics with high industrial relevance.
@@ III Methodology
We repeated this process three times in eleven months with roughly two-three months in between, depending on the amount of newly available videos. Repeating this process at three different time points allowed us to capture a more representative set of videos, as content on YouTube is continuously updated. This approach ensured that newer, potentially relevant videos published after the initial searches were included, improving the timeliness of our dataset.
@@
We transcribed the filtered videos using whisper-large [22] and Pyannote.audio for speaker diarization [23]. Next, we applied a semantic text splitter [24], to obtain disjunct, contextually coherent paragraphs of 300 tokens maximum, which corresponds to approximately 225 words, on average. This way, we yield 4100 paragraphs on which we applied topic modeling using BERTopic [25].
@@ IV Results
Next, we give an overview of our identified themes, including discussions on the individual aspects of each topic, for which we provide direct quotes. Each quote is set in gray color and includes a subscript identifier, linking it to the corresponding video in the references. We further investigate co-occurring topics in videos.
@@ IV-A Themes of Building LLM-based Applications
The most frequently mentioned topic across all themes is RAG [26]. While LLMs encode vast amounts of information, they are limited to the data available during training, missing out on newer information or proprietary, sensitive data. RAG addresses this limitation by leveraging a vector database to store and retrieve prior ingested data. Before generating a response, the database is queried for relevant documents that are subsequently added to the prompt. This does not only provide timely data, but also reduces hallucination by providing additional context to the prompt.
@@
Speakers explain some of the optimization challenges in greater detail, such as difficulties and trade-offs when determining a good chunk size. Chunking the data is usually done before ingestion in the database, when larger documents likely would not fit in the context window of the LLM or only parts of the document contain relevant information. However, chunking long documents might also lead to adverse cuts in context. One speaker explains this with an example of an article about Alan Turing, which would likely be not split correctly, such that the retrieval will [..] get me like the first three chunks, but the last three chunks, which also corresponds to the latter half of the document may not have a mention of Alan TuringV66. So, determining a good chunk size is crucial. One speaker advocates a human-in-the-loop process to determine good chunk sizes for different use cases: you’ll have a human who’s validating [..] this chunk size for this type of document is giving the best resultsV147. Newer approaches, such as contextual retrieval embed document context within individual chunks, but at the cost of larger chunk sizes.
The data within the vector database is stored as vectors, created by embedding models. How embeddings are created and updated is a frequent topic since the quality of your embeddings really drive the quality of a retrieval modelV15. However, pretrained embedding models might deliver an inadequate representation of your data [..].V32 and fine-tuning the embedding model is suited because [..] shifting the embedding points does help to improve performanceV32. Additionally, there is often the need to frequently update and extend data within the database to ensure there is no outdated information, which is a challenge on its own. Nonetheless the bigger problem is if the embedding models have to be exchanged or fine-tuned to account for new data, it may require re-indexing the whole database. This can become [..] very expensive and slow.V32. A potential solution lies in not fine-tuning the embedding model [..] but [..] fine tune some sort of adapter on the query and just keep the document embeddings [..] frozenV32. So, while RAG systems are seen as promising architectures, they pose their own challenges and trade-offs that need to be considered.
@@
A key factor for improving vanilla LLMs is to enhance the context length of the query, allowing the LLM to generate responses based on a larger quantity of information. However, the token size is usually fixed and it is up to the engineer to decide which information to store for later inclusion and which to drop. Thus, memory management within applications is essential to handle context efficiently. One speaker describes this as: At the end of the day, when I talk about memory in all our applications, what I really mean is remembering previous interactions and then using those to inform future interactionsV174.
Effective memory management involves selecting only the most relevant information to retain, rather than preserving all historical data, which ensures models remain efficient and avoid confusion from outdated or irrelevant context. Techniques such as conversation buffer memory and conversation summary memory have been proposed by the practitioners: the former keeps recent details accessible, while the latter provides a condensed summary of past interactions, balancing specificity and broader context. As another speaker notes, this method is an interesting way to address that problem for a chatbot-like scenarioV131 by focusing on the last $K$ relevant messages and managing context length effectively.
@@
A key insight shared was the importance of structured prompts, such as using XML tags or fixed prompt placement, to help models interpret context and improve performance. Additionally, certain prompt quirks—like using specific phrases or even statements like ”take a deep breath” or ”ensure the scores are correct”—sometimes enhance model responses, though the reasons remain unclear: So nobody knows why [..] you introduce like a statement, like ‘take a deep breath’ and the model tends to like do this.V66. This lack of theory makes prompting somewhat experimental, with practitioners trying unconventional statements based on observed improvements.
@@
Model size is not only a relevant for flexible deployment and low latency, but generally indicates the model’s capabilities and task performance. For many open ended tasks, a larger model size often leads to a higher quality of results, as it has been also shown via the observed scalability laws of LLMs [32]. Practitioners discuss this trade-off between model size and performance, alongside other relevant aspects, such as deployment efficiency. Currently, larger models exhibit better reasoning abilities but come with higher serving costs, requiring specific hardware and skills: [By using adaptation techniques] is how you’re able to build a much, much smaller model, which is equally as accurate, which is much, much easier to deploy.V96. Overall, this topic is closely related to specific techniques to improve model performance, such as fine-tuning or prompt engineering, as these technique generally hold the potential to tailor models towards a specific use case, and improve performance, especially for smaller models.
@@
In some cases, practitioners discuss how to optimize LLMs such that serving them is easier and less costly, while retaining performance on specific tasks. The techniques mentioned within discussions are quantization, knowledge distillation, and engineering optimizations. Quantization [33], for instance, reduces the size of the stored model weights to lower memory requirements and computations at the cost of reduced accuracy of the model. Knowledge distillation [34], in contrast, relies on training a smaller model as a student to replicate the behavior of the bigger model. Most parts of this topic correspond to explanations of these techniques, alongside pros and cons.
@@
There is lot of discussion about whether the LLM should be deployed in a dedicated cloud, on-premise, or whether using an external LLM API, such as provided by OpenAI [^2] https://openai.com/. This decision is not easy: speakers discuss several considerations, such as latency, performance, costs, and control. Hosting models on-premise or in a dedicated cloud has several advantages such as more predictable long term costs and better compliance and control, as noted by one speaker, [hosting your own model is cheaper] since you can control it, it’s a lot more predictableV16. Conversely, choosing LLMs-as-a-service is more convenient, has lower upfront costs and higher scalability. However, speakers mention several challenges they encountered with LLM APIs, such as rate limits, timeouts, and unpredictable downtimes, exemplary indicated by one speaker: [some days] you’ll just have really bad latency and your HTTP calls will time outV19. Another speaker mentions that they rely on a hybrid approach, using an external LLM API not for the latency sensitive use casesV17, but rather batch jobs that are not time sensitive. Hence, where to put the LLM is an outreaching architectural decision that needs to be made based on the tasks, resources, and environments at hand.
@@
Managing the infrastructure for training and inference of LLMs is a challenge on its own. One frequently mentioned tool is Kubernetes, a container orchestration platform that is already widely used for managing traditional ML workflows. One speaker, however, explicitly states that these workloads are not necessarily comparable: [..] a lot of machine learning 1.0 workloads were batch and I would say more the 2.0 like generative AI, most of that’s actually streamingV11. This possesses unique challenges for the kind of infrastructure that is needed and the tools to manage it. Thus, other hosting and optimization frameworks are necessary. One speaker mentions that they finally used Ray to spread the model across different instancesV103 as it was not possible to fit a model on a single Kubernetes node. Key considerations discussed are cost optimization and performance balancing that both are influenced by different techniques, such as load balancing, dynamic batching, and efficient resource allocation. LLM-specific techniques and patterns may evolve that have the potential to improve reliability and performance, such as described by one speaker: You need some form of distributed queue that is aware of the length of the completion and the batch sizes and that schedule is based on that.V51.
@@
Use case-specific evaluations with humans in the loop are often seen as gold standards, but are expensive and not scalable. An alternative is using LLMs to evaluate the output of another model, which is seen as very powerful as anything you can prompt a model to do, you can get it to evaluate.V3. While promising, this method also comes with its own challenges, such as ensuring that the evaluating model’s biases do not influence the assessment. Alongside different evaluation settings, speakers mention that having the right data and metrics is key for successful evaluation that builds trust throughout development cycle, from prototyping to monitoring and regression testing.
@@
A key challenge of using LLMs in an application is dealing with hallucinations, that is, when the model generates output that is factually incorrect or misleading. One speaker mentions that hallucinations might occur due to out-of-distribution requests, but [..] that quantifying what’s out of distribution is a lot more difficult [than in traditional ML].V4. Practitioners propose several strategies to mitigate hallucinations, such as prompt chaining and self-critique. While the former is about combining different prompts within one, the latter works by explicitly asking the model to reflect on its output: [..] you’re asking it to self critique, right? ‘Hey, does this make sense to you?’ And the model might say, ‘oh, I apologize. That actually doesn’t make sense’ and gets it right the second time.V3. Finally, one speaker indicates that by using LLMs, we can never fully exclude hallucinations and concludes: [..] well if my application cannot tolerate you know hallucinations maybe I shouldn’t use it for that application.V70.
@@
A recurring insight is that while larger context length is generally better since it allows LLMs to incorporate more in-context information and longer inputs, it is significantly more expensive as [..] when you take a bigger context window, [..] there is a quadratic increase in the cost of training.V159. At the same time, it is also noted that smaller context windows enable quicker experimentation during training and prompt engineering without excessive spending. So, these discussions emphasize the importance of finding an optimal context length, as it has substantial cost and performance implications for LLMs.
@@
Speakers discuss different tools, which work as an intermediary between the user and LLM and take care of the interactions by providing several mechanisms for handling LLM responses, ensuring correctness, safety, and compliance. Besides verifying outputs, these tools may include mechanisms to directly correct LLM output by automatically re-asking the model for correct or compliant output.